OpenStudy (anonymous):

3x+2y=-19 x-12y=19 solve using elimination

OpenStudy (whpalmer4):

@drewsorr1 They are supposed to use elimination, not substitution. @kkscout10 I started explaining the procedure to you when you first posted this. Why did you close and repost?

OpenStudy (cloverracer):

Can you insert this in the equation editor below please? :)

OpenStudy (anonymous):

@whpalmer4 You didn't respond

OpenStudy (whpalmer4):

I did. I asked you to do something. You didn't do it.

OpenStudy (anonymous):

@whpalmer4 Yes I did

OpenStudy (whpalmer4):

Go back and look. http://openstudy.com/users/kkscout10#/updates/534497f6e4b02bb88e659f8d

OpenStudy (texaschic101):

3x + 2y = -19 x - 12y = 19 --->(-3) -------------- 3x + 2y = -19 -3x + 36y = - 57 (result of multiplying by -3) --------------add 38y = - 76 y = -76/38 y = - 2 now sub -2 in for y in either equation to find x 3x + 2y = -19 3x + 2(-2) = -19 3x - 4 = -19 3x = -19 + 4 3x = - 15 x = -15/3 x = -5 check.. x - 12y = 19 -5 - 12(-2) = 19 -5 + 24 = 19 19 = 19 (correct) x = -5 and y = -2

OpenStudy (texaschic101):

do you have any questions ?

OpenStudy (anonymous):

No thanks I understand now